# ExoIris: fast exoplanet transmission spectroscopy in Python

### v1

@@ 2.7 Transit Timing Variations
Transit timing variations (deviations from strict periodicity in the times of transit centers, TTVs) are typically caused by gravitational interactions between planets in multi-planet systems. While TTVs are not an issue when analyzing transits individually, they must be accounted for in joint analyses of observations from multiple epochs. Failing to do so can distort the light curve shape and bias the inferred planet-star radius ratio, leading to inaccurate transmission spectra.
@@
2.8 Noise Modeling
@@
The white noise model assumes independent, normally distributed uncertainties, scaled by a single multiplicative factor per noise group. These uncertainties are provided as an array with the same dimensions as the flux array, containing one value per flux measurement. They can be supplied by the data reduction pipeline or estimated during dataset initialization. The noise scaling factor is included as a free parameter to account for possible under- or overestimation of the input uncertainties. This model is appropriate when the noise is dominated by photon noise and there is minimal time-correlated or systematic variability.
The GP model with fixed hyperparameters uses a user-defined kernel to describe correlated noise or long-term trends. Any kernel supported by celerite2 can be used, and ExoIris includes methods for optimizing hyperparameters for the Matern-3/2 kernel. This approach is suitable in most cases where correlated noise or longer-time-scale trends are present, and it offers a good balance between flexibility and computational efficiency.
The GP model with free hyperparameters allows the hyperparameters to be estimated jointly with the rest of the model parameters. Currently, this is supported only for the basic Matern-3/2 kernel and comes with a higher computational cost. Support for additional kernels may be added in future versions.
@@ 2.9 Masking of Bad Data
3 Example: WASP-39 b Early Release Science Joint Analysis
@@
The ExoIris example GitHub repository [^3] https://github.com/hpparvi/exoiris_examples contains a set of examples starting from simple analyses with one dataset and continuing to more complex cases combining many datasets into a joint analysis. The WASP-39 b Early Release Science (ERS) joint analysis provides a comprehensive demonstration of a joint transmission spectrum analysis using spectrophotometric light curves from WASP-39 b transits observed with NIRSpec, NIRSpec PRISM, NIRCam, and NIRISS. The data used in this example were produced by Carter et al. (2024), who reprocessed the original observations presented by Ahrer et al. (2023), Feinstein et al. (2023), Alderson et al. (2023), and Rustamkulov et al. (2023). The example illustrates how to configure bias, ephemeris, and noise groups for joint modeling and includes individual transmission spectroscopy analyses for each instrument.
3.1 Analysis Setup
Figure 5: ExoIris setup for the WASP-39 b ERS joint analysis. The top row shows the radius ratio knot locations, the second row shows the limb darkening knot locations, and the remaining rows show the central wavelengths of the spectroscopic elements for each dataset. The noise (N), offset (O), and epoch (E) group assignments are indicated beside each dataset’s wavelength coverage. Details on cropping and binning are provided in the text.
@@ 3.1 Analysis Setup
The joint analysis includes eight spectrophotometric datasets, as shown in Fig. 5. The NIRSpec PRISM data are cropped to $2.0\leq\lambda\leq 5.5$ $\mu\text{m}$ to remove saturated wavelengths (Carter et al., 2024; Rustamkulov et al., 2023), and then binned in time to 30 s (but not binned in wavelength). The NIRCam F322W2 data are cropped to $\lambda<4$ $\mu\text{m}$ and $t>2459783.32$, and binned in wavelength to a resolution of $R\approx 600$. The NIRSpec G395H data are split into two time segments ($2459791.45>t>2459791.62$ and $t>2459791.625$), cropped to $\lambda>2.85$ $\mu\text{m}$, and binned in wavelength to $R\approx 600$, resulting in four datasets. The NIRISS SOSS order 1 data are cropped to $0.86\leq\lambda\leq 2.8$ $\mu\text{m}$ and binned to $R\approx 300$, while order 2 data are cropped to $0.60\leq\lambda\leq 0.84$ $\mu\text{m}$ and binned to $R\approx 800$. Instead of using uncertainties from the reduction pipelines, we estimate a white noise scatter (standard deviation) for each light curve directly from the data.
@@
Each of the four transit observations is assigned to a separate epoch group, adding four transit center parameters to the model. Although WASP-39 b shows no signs of TTVs, the Carter et al. (2024) NIRISS reduction produces a significant offset in transit center time relative to a linear ephemeris, in contrast to the other datasets (see Fig. 10). While this offset is a data reduction artifact, it serves to demonstrate the functionality of epoch groups in ExoIris. Each detector is assigned a separate bias offset group, resulting in five free offset parameters. The NIRSpec PRISM dataset is used as the reference (offset group 0) due to its wide wavelength coverage. Each dataset is also assigned its own noise group, resulting in six white noise scaling parameters.
The transmission spectrum is modeled using 126 radius ratio knots across three wavelength regions: 0.6–0.75 $\mu\text{m}$ with a spacing of approximately 0.02 $\mu\text{m}$, a denser region around the 0.768 $\mu\text{m}$ K line with a spacing of 0.004 $\mu\text{m}$, and 0.78–5.5 $\mu\text{m}$ with a spacing of approximately 0.04 $\mu\text{m}$. Limb darkening is modeled with 10 knots linearly spaced from 0.6 to 5.5 $\mu\text{m}$.
3.2 Results
@@
Figure 7 shows a joint fit to the ”white” light curves. This is an optional initial step that can be used to initialize the transmission spectroscopy analysis, and is also useful for visual identification of spot crossings. Figure 8 shows the transmission spectrum estimated from the joint analysis, together with the transmission spectrum from Carter et al. (2024). Finally, Fig. 9 shows the residuals relative to the median posterior model, and Fig. 10 shows the posterior distributions for a set of quantities that are useful when assessing the quality of the observations and the analysis.
The ExoIris transmission spectrum agrees closely with the Carter et al. (2024) results with fitted limb darkening coefficients.
@@
As mentioned in Carter et al. (2024), JWST observations offer an unprecedented precision for transit-based stellar density estimation (assuming the eccentricity and argument of periastron are also known to a high precision.) Joint modeling of multiple spectrophotometric data sets naturally increases precision beyond that of individual analyses. Assuming a circular orbit, the stellar density precision for WASP-39 obtained from the joint model is 0.2%.
@@
Figure 7: WASP-39 b white light curves jointly fitted with ExoIris. Blue and yellow points show the observed photometry, and the black line shows the median posterior model. The NIRSpec NRS1 and NRS2 light curves are split in time due to a jump near mid-transit.
@@
Figure 8: WASP-39 b transmission spectrum estimated from the joint analysis of NIRSpec, NIRSpec PRISM, NIRCam, and NIRISS data using ExoIris. The black line shows the median of the posterior model, blue shading indicates the 16th and 84th percentiles, gray points show the radius ratio values from Carter et al. (2024), and the orange line shows the Carter et al. (2024) values binned to a similar resolution with the analysis.
Figure 9: Residuals relative to the median posterior model for the WASP-39 b joint analysis.
Figure 10: Posterior distributions from the WASP-39 b joint analysis for transit centers, bias offsets, white noise scaling factors, and out-of-transit baseline levels. Transit centers are shown relative to the best-fit linear ephemeris. The NIRISS dataset shows a significant discrepancy due to reduction artifacts in Carter et al. (2024), not physical TTVs. All datasets show small bias offsets relative to the PRISM data.
4 Discussion

### v2

@@ II.7 Transit Timing Variations
Transit timing variations (deviations from strict periodicity in the transit center times, TTVs) are typically caused by gravitational interactions between planets in multi-planet systems. While TTVs are not an issue when analyzing transits individually, they must be accounted for in joint analyses of observations from multiple epochs. Failing to do so can distort the light curve shape and bias the inferred planet-star radius ratio, leading to inaccurate transmission spectra.
@@
II.8 Spot Crossings and the Transit Light Source Effect
Stellar spot crossings and flux contamination from unocculted stellar spots and faculae—also known as the transit light source effect (TLSE; Pont et al., 2008; Sing et al., 2011; Pont et al., 2013; McCullough et al., 2014; Rackham et al., 2018, 2019)—all affect the inferred transmission spectra.
Individual stellar spot crossings are straightforward to identify, as these events produce characteristic bumps in the transit light curves. The TLSE arising from unocculted stellar inhomogeneities presents a more complex challenge. Both unocculted spots and faculae imprint wavelength-dependent signatures on the measured transit depths, with amplitudes and shapes determined by their temperatures and the stellar disk covering fractions.
ExoIris allows an arbitrary number of stellar spots to be included in the flux model. Each spot is associated with an epoch group and is modeled as a generalized Gaussian (Fig. 5),
$f(t,\lambda,T_{\star},T_{\mathrm{s}},c,a_{0},w,s)=\\ \delta(\lambda,T_{\star},T_{\mathrm{s}})a_{0}\exp\left(-\frac{2\;|t-c|}{w\ln(4)^{1/s}}\right)^{s},$ (10)
where $t$ is time, $\lambda$ is wavelength, $\delta$ is the wavelength-dependent spot contrast (normalized to unity at a reference wavelength $\lambda_{0}$), $T_{\star}$ and $T_{\mathrm{s}}$ are the photospheric and spot temperatures, respectively, $a_{0}$ is the spot amplitude at the reference wavelength, $w$ is the spot full width at half maximum, and $s$ is a shape parameter controlling the spot profile. The spot contrast, $\delta$, is computed from the stellar photosphere and spot temperatures using BT-Settl spectrum models (Allard et al., 2012).
Figure 5: Model used for stellar spot crossing events. The temporal shape of each event is modeled as a generalized Gaussian, while the wavelength dependence of the spot amplitude is computed using BT-Settl model spectra by Allard et al. (2012).
The effects from unocculted stellar heterogeneities can also be included into the joint model. Enabling TLSE modeling introduces the average temperatures of unocculted spots and faculae as free model parameters, and adds separate spot and faculae covering fractions as free parameters for each epoch group. The TLSE is then applied following the formalism by Rackham et al. (2018), where the transit model is multiplied by
$\epsilon_{\lambda,\mathrm{s+f}}=\frac{1}{1-f_{\mathrm{s}}\left(1-\frac{F_{\lambda,\mathrm{s}}}{F_{\lambda,\mathrm{p}}}\right)-f_{\mathrm{f}}\left(1-\frac{F_{\lambda,\mathrm{f}}}{F_{\lambda,\mathrm{p}}}\right)},$ (11)
where $f_{\mathrm{s}}$ and $f_{\mathrm{f}}$ denote the fractional disk coverage of spots and faculae for each epoch, and $F_{\lambda,\mathrm{p}}$, $F_{\lambda,\mathrm{s}}$, and $F_{\lambda,\mathrm{f}}$ are the fluxes from the stellar photosphere, spots, and faculae, respectively.
II.9 Noise Modeling
@@
The white noise model assumes independent, normally distributed uncertainties, scaled by a single multiplicative factor per noise group. These uncertainties are provided as an array with the same dimensions as the flux array, containing one value per flux measurement. They can be supplied by the data reduction pipeline or estimated during dataset initialization using the utility methods in ExoIris. The noise scaling factor is included as a free parameter to account for possible under- or overestimation of the input uncertainties. This model is appropriate when the noise is dominated by photon noise and there is minimal time-correlated or systematic variability. The log-likelihood is
$\log L=-\frac{1}{2}\sum_{j}^{n_{\mathrm{ds}}}\sum_{i}^{n_{\mathrm{p}(j)}}\left(\log\left(2\pi s_{\mathrm{g}(j)}^{2}\sigma_{j,i}^{2}\right)+\frac{\left(f_{j,i}-m_{j,i}\right)^{2}}{s_{\mathrm{g}(j)}^{2}\sigma_{j,i}^{2}}\right),$ (12)
where $n_{\mathrm{ds}}$ is the number of datasets, $n_{\mathrm{p}(j)}$ the number of points in each dataset (the two dimensions are here unrolled to one for simplicity), $s_{\mathrm{g}(j)}$ is the noise multiplier for the noise group assigned to the $j$th dataset, $f_{j,i}$ is the observed flux, $m_{j,i}$ is the model flux, and $\sigma_{j,i}$ is the flux uncertainty.
The log-likelihood for a Gaussian Process (GP) noise model is
$\log L=-\frac{1}{2}\sum_{j}^{n_{\mathrm{ds}}}\left(n_{\mathrm{p}(j)}\log(2\pi)+\log\left|\boldsymbol{\Sigma}_{j}\right|+\mathbf{r}_{j}^{\mathsf{T}}\boldsymbol{\Sigma}_{j}^{-1}\mathbf{r}_{j}\right),$ (13)
where $\mathbf{r}_{j}=\mathbf{f}_{j}-\mathbf{m}_{j}$ is the residual vector containing the observed fluxes $\mathbf{f}_{j}$ and model fluxes $\mathbf{m}_{j}$. The covariance matrix, $\boldsymbol{\Sigma}$, is constructed as
$\boldsymbol{\Sigma}=\mathbf{K(t,t)}+\boldsymbol{\sigma}^{2}\mathbf{I},$ (14)
where $\mathbf{K}$ is the covariance kernel with either fixed or fitted hyperparameters and $\mathbf{t}$ are the mid-exposure times. The GP noise model considers only time-dependent systematics.
The GP model with fixed hyperparameters uses a user-defined kernel to describe correlated noise or long-term trends. The hyperparameters can be defined separately for each noise group, but they cannot vary along the wavelength or time axes. Any kernel supported by celerite2 can be used, and ExoIris includes methods for optimizing hyperparameters for the Matérn-3/2 kernel. This approach is suitable in most cases where correlated noise or longer-time-scale trends are present, and it offers a good balance between flexibility and computational efficiency.
The GP model with free hyperparameters allows the hyperparameters to be estimated jointly with the rest of the model parameters. Currently, this is supported only for the basic Matérn-3/2 kernel and comes with a higher computational cost. Support for additional kernels may be added in future versions.
The GP is evaluated independently for each wavelength channel using celerite2. Internally, this is implemented by reshaping each two-dimensional residual array of size $(n_{\mathrm{wl}},n_{\mathrm{time}})$ into a one-dimensional array of length $n_{\mathrm{wl}}\times n_{\mathrm{time}}$. The corresponding time array is repeated $n_{\mathrm{wl}}$ times, and a constant time offset is added between successive repetitions such that different passbands are separated by a sufficiently large temporal gap. This ensures that the covariance between points belonging to different passbands is effectively zero. The GP log-likelihood for all passbands is then evaluated in a single call using the combined time array and the flattened residuals.
@@ II.10 Masking of Bad Data
II.11 Optimization and Posterior Inference
The transmission spectrum inference in ExoIris is divided into a global optimization phase and a subsequent posterior sampling phase. The first phase uses a Differential Evolution (DE) optimizer (Storn and Price, 1997; Price et al., 2005) to identify the region of the global posterior mode. The optimization is initialized with a population of $n$ parameter vectors, and, during the optimization, the DE algorithm concentrates the parameter vector population toward the global posterior mode.
Transition to the sampling phase occurs once the population has concentrated into a sufficiently tight clump. Convergence is determined by the width of the log-posterior distribution within the population, and the convergence limit is user-adjustable. Posterior inference is then performed using the affine invariant Markov Chain Monte Carlo (MCMC) sampler implemented in the emcee package (Foreman-Mackey et al., 2013). By initializing the sampler with the clumped population from the optimization phase, the Markov chains start their exploration near the global posterior mode.
Both optimization and sampling support parallel execution via Python’s multiprocessing module or the mpi4py package. Additionally, both optimization and MCMC sampling can be performed semi-interactively. The optimizer can be executed for a specific number of iterations, after which the current best-fit model and population state can be examined using various diagnostics. The optimization can then be resumed for additional iterations, continuing directly from its previous state without loss of progress. Similarly, the MCMC sampler can be run for a set number of iterations to permit the inspection of intermediate results, such as chain convergence and autocorrelation times, before continuing sampling from the last recorded state.
The state of an ExoIris analysis can be saved at any time. Every saved model is fully self-contained because the save file stores all information required to reproduce the analysis, including datasets, priors, and the current state of the parameter population. This architecture enables iterative and branching workflows in which a saved analysis serves as the basis for subsequent analyses. Users can introduce additional model complexity, such as adding radius-ratio knots, freeing knot locations, incorporating the transit-light-source effect, or updating datasets, while leveraging the progress made in earlier stages of the analysis.
II.12 Atmosphere Retrieval with Oversampled Spectra
After sampling, the ExoIris.transmission_spectrum utility method can be used to generate a posterior transmission spectrum. This method computes the mean transmission spectrum, $\boldsymbol{\mu}$, and its covariance matrix, $\boldsymbol{\Sigma}$, on an arbitrary wavelength grid with $n_{\lambda}$ elements. Additionally, the ExoIris.create_loglikelihood_function method can be used to create a callable LogLikelihood object ready for use in atmospheric retrieval frameworks.
A specialized approach to calculating the log-likelihood is required because the transmission spectrum is oversampled. That is, the transmission spectrum is derived by interpolating the posterior distribution of $n_{\mathrm{k}}$ radius-ratio knots, where, typically, $n_{\mathrm{k}}\ll n_{\lambda}$, and the transmission spectrum elements in $\boldsymbol{\mu}$ are strongly correlated across wavelengths. Consequently, the points cannot be treated as independent observations, and the full covariance structure must be accounted for when evaluating the likelihood of theoretical atmospheric models.
Ideally, a theoretical transmission spectrum $\mathbf{m}$ would be compared to the posterior using the standard multivariate normal log-likelihood,
$\log L=-\frac{1}{2}\left(n_{\lambda}\log(2\pi)+\log(|\boldsymbol{\Sigma}|)+\mathbf{r}^{\mathsf{T}}\boldsymbol{\Sigma}^{-1}\mathbf{r}\right),$ (15)
where $\mathbf{r}=\boldsymbol{\mu}-\mathbf{m}$ is the residual vector. However, the interpolation process renders the posterior transmission spectrum rank-deficient. For linear interpolation methods (nearest-neighbor, linear, and B-spline), the rank of $\boldsymbol{\Sigma}$ is strictly limited to $n_{\mathrm{k}}$ instead of $n_{\lambda}$ because the spectrum at any wavelength is a linear combination of the $n_{\mathrm{k}}$ knot values. For non-linear interpolators (PCHIP, Makima), while the algebraic rank may exceed $n_{\mathrm{k}}$, the effective rank still remains constrained to $\approx n_{\mathrm{k}}$. This rank deficiency implies that $\boldsymbol{\Sigma}$ is ill-conditioned and that its direct inversion is numerically unstable (or plain impossible), rendering the standard likelihood evaluation infeasible.
To construct a robust likelihood, ExoIris adopts the subspace compression formalism, often referred to as Karhunen-Loève (KL) compression, widely used in cosmology to handle rank-deficient datasets (e.g., Tegmark et al., 1997). The code performs an eigendecomposition of the empirical covariance,
$\boldsymbol{\Sigma}=\mathbf{U}\,\boldsymbol{\Lambda}\,\mathbf{U}^{\mathsf{T}},$ (16)
and identifies the eigenvalues that are significantly larger than numerical noise. Only these eigenvalues and their corresponding eigenvectors represent directions in which the posterior distribution has meaningful support. Eigenvalues smaller than a given tolerance are discarded. The retained eigenvectors, $\mathbf{U}_{r}$, form an orthonormal basis for the reduced subspace, and the retained eigenvalues (variances along the eigenvectors) form the diagonal matrix $\boldsymbol{\Lambda}_{r}$.
The likelihood is then evaluated by projecting the residuals onto the subspace defined by the eigenvectors. The resulting subspace log-likelihood is
$\log L=-\frac{1}{2}\left(K\log(2\pi)+\sum_{i=1}^{K}\log\lambda_{i}+\sum_{i=1}^{K}\frac{(\mathbf{u}_{i}^{\mathsf{T}}\mathbf{r})^{2}}{\lambda_{i}}\right),$ (17)
where $\mathbf{u}_{i}$ is the $i$th eigenvector, $\lambda_{i}$ is the $i$th eigenvalue, and the summation runs only over the $K$ valid eigenvectors.
III Example: WASP-39 b Early Release Science Joint Analysis
@@
The ExoIris example GitHub repository [^3] https://github.com/hpparvi/exoiris_examples contains a set of examples ranging from simple analyses with a single dataset to more complex cases combining multiple datasets into a joint analysis. The WASP-39 b Early Release Science (ERS) joint analysis provides a comprehensive demonstration of a joint transmission spectrum analysis using spectrophotometric light curves from WASP-39 b transits observed with NIRSpec, NIRSpec PRISM, NIRCam, and NIRISS. The data used in this example were produced by Carter et al. (2024), who gathered the original observations presented by Ahrer et al. (2023), Feinstein et al. (2023), Alderson et al. (2023), and Rustamkulov et al. (2023). The example illustrates how to configure bias, ephemeris, and noise groups for joint modeling and includes individual transmission spectroscopy analyses for each instrument.
III.1 Analysis Setup
@@
Figure 7: ExoIris setup for the WASP-39 b ERS joint analysis. The top row shows the radius-ratio knot locations, the second row shows the limb-darkening knot locations, and the remaining rows show the central wavelengths of the spectroscopic elements for each dataset. The noise (N), offset (O), and epoch (E) group assignments are indicated beside each dataset’s wavelength coverage. Details on cropping and binning are provided in the text.
@@
The joint analysis includes eight spectrophotometric datasets shown in Fig. 6. The NIRSpec PRISM data are cropped to $2.0\leq\lambda\leq 5.5$ $\mu\text{m}$ to remove saturated wavelengths (Carter et al., 2024; Rustamkulov et al., 2023), and then binned in time to 30 s (but not binned in wavelength). The NIRCam F322W2 data are cropped to $\lambda<4$ $\mu\text{m}$ and $t>2459783.32$, and binned in wavelength to a resolution of $R\approx 600$. The NIRSpec G395H data are split into two time segments ($2459791.45<t<2459791.62$ and $t>2459791.625$), cropped to $\lambda>2.85$ $\mu\text{m}$, and binned in wavelength to $R\approx 600$, resulting in four datasets. The NIRISS SOSS order 1 data are cropped to $0.86\leq\lambda\leq 2.8$ $\mu\text{m}$ and binned to $R\approx 300$, while order 2 data are cropped to $0.60\leq\lambda\leq 0.84$ $\mu\text{m}$ and binned to $R\approx 800$. Instead of using uncertainties from the reduction pipelines, I estimate the white-noise scatter (standard deviation) for each light curve directly from the data.
@@
Each of the four transit observations is assigned to a separate epoch group, adding four transit center parameters to the model. Although WASP-39 b shows no signs of TTVs, the Carter et al. (2024) NIRISS reduction produces a significant offset in transit center time relative to a linear ephemeris, in contrast to the other datasets. While this transit time offset is a data-reduction artifact, it demonstrates the functionality of epoch groups in ExoIris. Each detector (and NIRISS SOSS order) is assigned a separate bias offset group, resulting in five free offset parameters. The NIRSpec NRS1 dataset is used as the reference (offset group 0). Each dataset is also assigned to its own noise group, resulting in six white-noise scaling parameters. The final model setup is illustrated in Fig. 7.
The transmission spectrum estimation is carried out progressively, starting with a model that uses nearest-neighbor radius-ratio interpolation and 25 radius-ratio knots, $n_{\mathrm{k}}$. After fitting and sampling the initial model, the model resolution is increased to $n_{\mathrm{k}}=60$, $n_{\mathrm{k}}=130$, and $n_{\mathrm{k}}=230$, every time saving the models and the posterior sampling results separately. The region around the 0.768 $\mu\text{m}$ K line is sampled with a higher radius-ratio knot resolution after the initial run. The analyses are repeated for the linear, quadratic B-spline, cubic B-spline, PCHIP, and Makima radius-ratio interpolation schemes. Limb darkening is modeled using 10 linearly spaced knots from 0.6 to 5.5 $\mu\text{m}$. In the initial low-resolution analysis, the limb-darkening coefficients are constrained with loose, informative priors derived using LDTk, whereas in subsequent analyses the priors are changed to uninformative uniform priors.
III.2 Results
@@
Figure 8 shows a joint fit to the white-light curves. This is an optional initial step that can be used to initialize the transmission spectroscopy analysis, and is also useful for visual identification of spot crossings. Figure 9 shows the WASP-39 b transmission spectrum estimates from four progressively complex joint analyses, and Fig. 10 compares the ExoIris-estimated transmission spectrum with quadratic B-spline radius-ratio interpolation and $n_{\mathrm{k}}=130$ against the results by Carter et al. (2024). Finally, Fig. 11 shows the posterior distributions for a set of ExoIris model parameters.
@@
The ExoIris transmission spectrum agrees well with the Carter et al. (2024) spectra. The differences in the NIRSpec PRISM and NIRCam mean levels (Fig. 10) are due to the instrument-dependent offsets discussed in Carter et al. (2024). These offsets are captured by the ExoIris analysis, as shown in Fig. 11. Additionally, as mentioned in Carter et al. (2024), JWST observations offer unprecedented precision for transit-based stellar density estimation (assuming the eccentricity and argument of periastron are also known to high precision). Joint modeling of multiple spectrophotometric datasets naturally increases precision beyond that of individual analyses. Assuming a circular orbit, the ExoIris analysis yields a stellar density of $1.648\pm 0.002$ and an impact parameter of $0.4679\pm 0.0006$.
@@
Figure 8: WASP-39 b white-light curves jointly fitted with ExoIris. Blue and yellow points show the observed photometry, and the black line shows the median posterior model. The NIRSpec NRS1 and NRS2 light curves are split in time due to a jump near mid-transit.
@@
Figure 9: Posterior WASP-39 b transmission spectra estimated from the joint analysis of NIRSpec, NIRSpec PRISM, NIRCam, and NIRISS data using ExoIris. The black line shows the posterior median model, and the blue shading shows the 68% and 95% central posterior intervals. The transmission spectrum was estimated progressively, starting with nearest-neighbor interpolation with a low radius-ratio knot resolution (top), and ending with a linear interpolation with a high knot resolution (bottom).
Figure 10: As in Fig. 9, but comparing the Carter et al. (2024) results with the ExoIris analysis using quadratic B-spline interpolation with 130 radius-ratio knots. The black points show the median posterior radius-ratio knot values, and the orange shading shows the $1\sigma$ intervals from the Carter et al. (2024) analysis with fitted limb darkening and bin scale 5.
Figure 11: Posterior distributions for a subset of ExoIris model parameters from the WASP-39 b joint analysis. The first panel from the top shows the transit center time offsets for the four epoch groups, the second panel shows the additive offset posteriors, and the two bottom panels compare the stellar density and impact parameter posteriors to the estimates presented in Carter et al. (2024).
IV Discussion